OpenStudy (anonymous):

AB passes through A(-3, 0) and B(-6, 5). What is the equation of the line that passes through the origin and is parallel to AB ? A) 5x – 3y = 0 B) -x + 3y = 0 C) -5x – 3y = 0 D) 3x + 5y = 0 E) -3x + 5y = 0

OpenStudy (anonymous):

@SolomonZelman can you please help me!!!(:

OpenStudy (solomonzelman):

Okay, first you have to find the slope of the line that passes through (the given points) A and B. Can you do that ?

OpenStudy (anonymous):

is it 5/-3 ?

OpenStudy (solomonzelman):

YES \(\LARGE\color{blue}{ \rm \frac{(0)-(5)}{(-3)-(-6)}=\frac{0-5}{-3+6} =\frac{-5}{3} }\)

OpenStudy (solomonzelman):

So, now we can say \(\LARGE\color{blue}{ \rm y=-\frac{5}{3}~x~+b }\) And when the line goes through the origin, the "b" (the y-intercept) is zero.

OpenStudy (anonymous):

soo it would be C?

OpenStudy (solomonzelman):

Yes.

OpenStudy (anonymous):

thank you so much!!